OpenStudy (anonymous):

Y varies inversely with x and y = -12 when x = 6. A. xy = -2 B. xy = -72 C. xy = -0.5

OpenStudy (anonymous):

xy = -72

OpenStudy (anonymous):

\[y \alpha \frac{1}{x}\] \[y = \frac{k}{x}\] find K by subsitituiting the values of x and y in this equation k=12(-6) = 72